8. A student counted the number of snails at a pond in east Edmonton in April and found 38 snails. Four months later, the stu can someone help me with this whole page questions. i dont know how to do it. please do it right and thank you
Answer #1

Ans 8) population of snail in April = 38 = initial population = N0

Population of snail after 4 months = 214 = final population = Nt

Time = t = 4 months = 4/12 years = 1/3 year

Ans a) growth rate of snail population.

R = ∆N / t

R = growth rate

∆N = Nt - N0

t = time

Putting all values in formula. We get,

R = 214 - 38 / 1/3.

= 176 * 3/1

= 528 snails/year

Ans b) per capita growth rate of population is

r = (∆N/N0 * 100)/t

r = (176/38 * 100) / 1/3

= (17600/38) / 1/3

= 463.16 / 1/3

= 463.16 * 3

= 1389.48% per year

Ans 9) 2 factors that determine where population is distributed are :

1) physical factors ---:> include the physical conditions of a place like the climatic conditions, soil type, landform, water bodies,etc. These are the basic requirements of population to live in.

2) human factors -----> include the human nature or behaviour like the culture and tradition, the economic conditions, politics, history.

Ans 10) the population distribution pattern in Canada is not uniform. Most people live in southern part than in Northern part.